Prove that the equation of a line through (π‘₯1 , 𝑦1) and (π‘₯2 , 𝑦2) can be expressed in the form determinant

x y 1

x1 y1 1

x2 y2 1

= 0


1
Expert's answer
2022-02-08T11:36:24-0500

Solution:

Two points are (π‘₯1 , 𝑦1) and (π‘₯2 , 𝑦2).

Slope "=m=\dfrac{y_2-y_1}{x_2-x_1}"

Intercept"=c=\dfrac{x_1y_2-x_2y_1}{x_2-x_1}"

Now, equation is "y=mx+c"

"\Rightarrow y=\dfrac{y_2-y_1}{x_2-x_1}x+\dfrac{x_1y_2-x_2y_1}{x_2-x_1}\ ...(i)"

Now, given determinant is "\begin{vmatrix} x&y&1\\x_1&y_1&1\\x_2&y_2&1\end {vmatrix}=0"

On expanding,

"x(y_1-y_2)-y(x_1-x_2)+(x_1y_2-x_2y_1)=0
\\\Rightarrow x\dfrac{(y_1-y_2)}{(x_1-x_2)}-y+\dfrac{(x_1y_2-x_2y_1)}{(x_1-x_2)}=0
\\ \Rightarrow y=\dfrac{y_2-y_1}{x_2-x_1}x+\dfrac{x_1y_2-x_2y_1}{x_2-x_1}\ ...(ii)"

From (i) & (ii), it is proved.